# Breaking the HBM Bit Cost Barrier: Domain-Specific ECC for AI Inference Infrastructure

### v1

@@ IV-C Importance-Adaptive Protection
Fig. 8 presents the impact of targeted bit flips in different BF16 fields on model accuracy for the LLaMA 3.1 8B model. We randomly inject faults into the sign, exponent, and mantissa fields by flipping bits at rates ranging from $10^{-8}$ to $10^{-3}$. The model is evaluated on two benchmarks: PIQA [3], which assesses physical reasoning, and MMLU [4]. The results reveal that sign and mantissa corruption has minimal effect, even at the highest error rate. Accuracy remains above 97% of the error-free baseline, demonstrating high resilience to noise in these bit-planes. In contrast, exponent errors result in severe degradation: at a flip rate of $10^{-3}$, PIQA accuracy drops to 61.7% and MMLU to 38.1%. These findings indicate that the exponent bits are far more critical for preserving numerical integrity, and thus should be prioritized in error protection schemes.
Figure 8: Normalized LLaMA 3.1 8B accuracy under bit flips in BF16 sign, exponent, and mantissa fields on (a) PIQA and (b) MMLU.
Fig. 9 explores the efficiency gains of concentrating ECC protection solely on the exponent field. We compare effective HBM bandwidth utilization between exponent-only and full-bit ECC strategies across RS codeword lengths ranging from 64B to 2048B, under three increasingly aggressive BERs ($10^{-5}$, $10^{-4}$, and $10^{-3}$). At $10^{-5}$, exponent-only ECC achieves 98.43% utilization at 64B and 98.49% at 2048B, while full-bit ECC reaches only 93.91% and 94.73%, respectively. As the BER rises to $10^{-4}$, the efficiency gap widens further. For instance, with 256B codewords, exponent-only ECC maintains 98.23% bandwidth, compared to 90.39% for full-bit ECC. Even under extreme noise at $10^{-3}$, exponent-only protection holds up: 93.48% at 64B and 95.69% at 2048B, whereas full-bit ECC drops to 80.90% and 89.00%, respectively. Across all settings, adaptive protection consistently reduces ECC-related traffic overhead, with bandwidth improvements reaching up to 12.6%.
Together, these results demonstrate the effectiveness of importance-aware ECC in balancing robustness and efficiency. By aligning protection strength with data sensitivity, focusing on the exponent field, this approach reduces bandwidth consumption and decoder workload, while preserving accuracy under high error rates. Importance-adaptive ECC thus offers a compelling direction for lowering HBM cost and controller complexity, opening the door to more scalable and resilient AI memory systems.
Figure 9: Effective HBM bandwidth utilization for exponent-only and full-bit ECC under BERs of $10^{-5}$, $10^{-4}$, and $10^{-3}$ under different codeword lengths.

### v2

@@ IV-C Importance-Adaptive Protection
Our importance-adaptive protection scheme is inspired by the motivational study in Fig. 7, which reports normalized accuracy under targeted BF16 bit flips for LLaMA-3.1-8B [9], Voxtral-Mini-3B [10], and Qwen3-4B [11] on PIQA [12] and MMLU [13] as we sweep flip rates from $10^{-8}$ to $10^{-3}$. This is a stress test on unprotected bit-planes to rank vulnerability. The trend is consistent: exponent flips are the dominant failure mode, while sign and mantissa flips have much smaller impact. At $10^{-3}$ with exponent-only corruption, PIQA drops to a worst case of 61.7% and MMLU to 38.1%. In contrast, if we protect the exponent and allow faults only in sign or mantissa, accuracy remains high even at $10^{-3}$ in the worst case: on PIQA it is at least 97.3% (sign) and 96.8% (mantissa). On MMLU it is at least 84.2% (sign) and 94.0% (mantissa). These results motivate an importance-adaptive policy that always protects exponent bit-planes and uses reduced or no protection for sign/mantissa, cutting ECC load with minimal accuracy loss. We show BF16 here for space, the mechanism is format-agnostic.
Figure 7: Motivational study demonstrating the critical vulnerability of exponent bits in unprotected BF16 models. Accuracy is evaluated on PIQA (top) and MMLU (bottom) after injecting bit-flips into the sign, exponent, and mantissa.
Fig. 8 explores the efficiency gains of concentrating ECC protection solely on the exponent field (on LLaMA-3.1-8B and MMLU). We compare effective HBM bandwidth utilization between exponent-only and full-bit ECC strategies across RS codeword lengths ranging from 64B to 2048B, under three increasingly aggressive BERs ($10^{-5}$, $10^{-4}$, and $10^{-3}$). At $10^{-5}$, exponent-only ECC achieves 98.4% utilization at 64B and 98.5% at 2048B, while full-bit ECC reaches only 93.9% and 94.7%, respectively. As the BER rises to $10^{-4}$, the efficiency gap widens further. For instance, with 256B codewords, exponent-only ECC maintains 98.23% bandwidth, compared to 90.4% for full-bit ECC. Even under extreme noise at $10^{-3}$, exponent-only protection holds up: 93.5% at 64B and 95.7% at 2048B, whereas full-bit ECC drops to 80.9% and 89.0%, respectively. Across all settings, adaptive protection consistently reduces ECC-related traffic overhead, with bandwidth improvements reaching up to 12.6%.
Therefore, in practice, we always protect the exponent planes. Together these results demonstrate the effectiveness of importance-aware ECC in balancing robustness and efficiency. By aligning protection strength with data sensitivity, focusing on the exponent field, this approach reduces bandwidth consumption and decoder workload, while preserving accuracy under high error rates. Importance-adaptive ECC thus offers a compelling direction for lowering HBM cost and controller complexity, opening the door to more scalable and resilient AI memory systems.
Figure 8: Effective HBM bandwidth utilization on LLaMA-3.1-8B (MMLU) for exponent-only vs. full-bit ECC under BERs of $10^{-5}$, $10^{-4}$, and $10^{-3}$ across codeword lengths.